OpenStudy (anonymous):

(-7x^4y)(-10x^6y^2) Multiply. (^ stands for the power)

OpenStudy (unklerhaukus):

\[\large (-7x^4y)(-10x^6y^2)\\ \large=(-7\times-10)(x^4\times x^6)(y\times y^2)\\ =\]

OpenStudy (unklerhaukus):

\[x^4\times x^6=x^{4+6}\]

OpenStudy (anonymous):

So would it be 70x^10y^3?

OpenStudy (unklerhaukus):

\[\checkmark\]

OpenStudy (anonymous):

Thank you!

OpenStudy (unklerhaukus):

\[\ddot\smile\]